OpenStudy (anonymous):

I am stuck, I am trying to solve trigonometric equation. sin (x/2) = -1/2, All real numbers

OpenStudy (phi):

do you know the graph for sin( A ) ? find all the A's where sinA = -½ then A= x/2 so x= 2A

OpenStudy (anonymous):

\[\frac{ 7\Pi }{ 6 }\] and \[\frac{ 11\Pi }{ 6 }\] equal sin(a)= -\[\frac{ 1 }{ 2 }\]

OpenStudy (anonymous):

then would i just solve for x?

OpenStudy (anonymous):

@phi

OpenStudy (phi):

x has to be twice those angles (so that when you find x/2, you get those angles, and sin()= -½)

OpenStudy (anonymous):

ok meaning my answer would be \[\frac{ 7\Pi }{ 3 } +4\Pi n\]?

OpenStudy (phi):

that looks good also 11 pi/3 + 4 pi n

OpenStudy (anonymous):

Do you think i should always add the other possible answer?

OpenStudy (anonymous):

thank you so much ^.^

OpenStudy (phi):

yes, unless they end up being the same numbers. (I have not checked... they might be)

OpenStudy (phi):

they are different, so you should include both

OpenStudy (anonymous):

Ok thank you ^.^